OpenStudy (anonymous):

Solve. a + 1 = √b+1 for b

OpenStudy (anonymous):

A. b = a^2 + 2a B. b = a + 2a^2 C. b = 2a^2 + a

OpenStudy (studygurl14):

Hello @vegangirl063 ! Welcome to Openstudy! Is it like this \(\Large a+1=\sqrt{b}+1\) or like this? \(\Large a+1=\sqrt{b+1}\)

OpenStudy (anonymous):

Thank you! =)) and its the second one

OpenStudy (studygurl14):

Okay. So we're solving for b. That means we need to get b alone on one side of the = sign. Notice how b is under a sqrt? Do you know how to get it out? Hint: What's the opposite of taking the sqrt of something?

OpenStudy (bee_see):

cancel the square root by multiplying by the 2nd power on both sides.

OpenStudy (bee_see):